OpenStudy (anonymous):

a parabola has a vertex with coordinates -4,1 and goes through the points coordinates 2,-17. find the equation of the parabola in the form of y=ax^2+bx+c?

OpenStudy (jdoe0001):

hmm do you know the vertex form of a parabola?

OpenStudy (anonymous):

some of it i dont unde stand this question if u an explain it im sue ill undestand

OpenStudy (jdoe0001):

well if you notice at the figure http://www.mathwarehouse.com/geometry/parabola/images/standard-vertex-forms.png the one on the right side is the vertex form so we know where the vertex is at, (-4, 1) that means the (h, k) are those values h = -4 k= 1 thus \(\bf y = a(x-h)^2+k \implies y = a(x-(-4))^2+ 1\implies y = a(x+4)^2+ 1\)

OpenStudy (jdoe0001):

so we just need "a" so.. what's "a" anyway? well we dunno but we know that another point, besides the vertex, in the parabola is (2, -17), since it goes through it so let us use that ( 2, -17) meaning x = 2 y = -17 thus \(\bf y = a(x-h)^2+k \implies y = a(x-(-4))^2+ 1\implies y = a(x+4)^2+ 1\\ \quad \\ (2\quad ,\quad -17)\implies (-17) = a((2)+4)^2+ 1\) solve for "a" to get "a"

OpenStudy (jdoe0001):

once you get "a', to get the standard form just expand the binomial

OpenStudy (anonymous):

k thanks for help

OpenStudy (jdoe0001):

yw